A jet lands on a aircraft carrier at 750 mile (100 m/s).
(a)what is its acceleration if it stops in 10 sec.
(b)what is the displacement of the plane which it is stopping.
1
Expert's answer
2020-03-02T09:09:45-0500

a)


"a=\frac{v}{t}=\frac{100}{10}=10\frac{m}{s^2}"

b)


"s=0.5at^2=0.5vt"

"s=0.5(100)(10)=500\ m"